# FCRF: Flexible Constructivism Reflection for Long-Horizon Robotic Task Planning with Large Language Models

### v1

@@ Abstract
Autonomous error correction is critical for domestic robots to achieve reliable execution of complex long-horizon tasks. Prior work has explored self-reflection in Large Language Models (LLMs) for task planning error correction; however, existing methods are constrained by inflexible self-reflection mechanisms that limit their effectiveness. Motivated by these limitations and inspired by human cognitive adaptation, we propose the Flexible Constructivism Reflection Framework (FCRF), a novel Mentor-Actor architecture that enables LLMs to perform flexible self-reflection based on task difficulty, while constructively integrating historical valuable experience with failure lessons. We evaluated FCRF on diverse domestic tasks through simulation in AlfWorld and physical deployment in the real-world environment. Experimental results demonstrate that FCRF significantly improves overall performance and self-reflection flexibility in complex long-horizon robotic tasks. Website at https://mongoosesyf.github.io/FCRF.github.io/
@@ III-B Self-Reflection Process in LLMs Planning
Discussing LLMs task planning process with self-reflection and memory module $mem$, the self-reflection content for current unsuccessful planning trail $sr^{t}\in SR$ is generated by the LLM for self-reflection $M_{sr}$, while the reflection process can be described as $M_{sr}(sr^{t}|s^{t},mem^{t})$, meaning that the self-reflection content $sr^{t}$ of the current trail $t$ is generated by $M_{sr}$ based on the current state $s^{t}$, and persistent memory $mem^{t}$ lasts until the current trail including existing task trajectory.
The self-reflection generated by $M_{sr}$ would be stored in the agent memory $mem$ and passed to planner LLM in the next trail, to generate better action sequence $\{a^{t+1}_{1},\dots,a^{t+1}_{i}\}$ as the new output of the planning task. The sequence of actions is generated by the strategy function $\Phi(a^{t+1}_{1},\dots,a^{t+1}_{i}|G,s^{t+1},o^{t+1},sr^{t})$, meaning that new action sequence $\{a^{t+1}_{1},\dots,a^{t+1}_{i}\}$ from $step_{1}$ to $step_{i}$ is generated based on the task goal $G$, the current agent state $s^{t+1}$, the current observation of task environment $o^{t+1}$, and the generated self-reflection content $sr^{t}$ of last failed trail.
@@ IV-A LLM as Actor
where $A^{t}=\{a^{t}_{1},\dots,a^{t}_{i}\}$ denotes the current action sequence, $G$ represents the task goal, $o^{t}$ represents the current environmental observations during trail $t$, $SR^{t-1}=\{sr^{t-k},\dots,sr^{t-1}\}$ represents the set of self-reflection among the past $k$ trials to consider about, and $\theta$ represents the parameters of $M_{a}$. The action sequence generated by Actor LLM $M_{a}$ would be stored in the memory module $mem$ for the following self-reflection process of the task, until the current task is successfully completed.
@@ IV-B LLM as Constructivism Self-Reflection Mentor
The constructivism self-reflection process of our framework is guided by an LLM prompted as a $Mentor$, represented as $M_{m}$, which contains three sub-modules $M_{exp}$, $M_{lesson}$ and $M_{cons}$. The reflection process is divided into three parts: the summary of valuable experience, the summary of failure lessons, and the comprehensive construction of the two above parts. The final generated reflection content will be added to the memory module $mem$ and can be called by the Actor LLM $M_{a}$ in subsequent trails for planning.
@@
After completing the valuable experience summary process and the failure lesson summary process, the constructor submodule $M_{cons}$ will execute the comprehensive construction process to comprehensively construct the successful and failed experience, integrate the final reflection result in the form of an improved plan. During the construction process, $M_{cons}$ integrates the summarized valuable experience $Exp^{t}$ with the newly injected $Lesson^{t}$, unifying them into a sequence of actions, which forms the new constructive plan to be executed in the next trail. The reflected new constructive plan will be stored in the $mem$ module and guide the planning process of $M_{actor}$ in the next trail. The comprehensive construction process can be described as:
@@ IV-C Memory Management
During the task planning process, it is important for LLMs to maintain context memory. However, if all memory contents are included in prompts, it will lead to redundant prompts, increasing the memory burden of LLMs, even affect their performance. Motivated by the long-short-term memory mechanism of previous work Reflexion, we design the memory module $mem$ for comprehensive memory management, which is equivalent to an external buffer of LLMs. The $mem$ is divided into trajectory management module $Traj^{t}$ and reflection management module $Refl^{t}$, comprehensively manages the memory of task trajectories and reflection contents. The $Traj^{t}$ module is a short-term memory module storing the task trajectory $\tau^{t}$ of the current trail $t$. The $Refl^{t}$ module is a long-term memory module storing all failure self-reflection contents up to trail $t$, which can be represented as $\{sr^{1},\dots,sr^{t}\}$. The reflection contents are stored for targeted reading and calling by the Mentor LLM $M_{m}$ in the following trails. Our memory mechanism $mem$ relieves the memory pressure of LLMs, improving the efficiency of memory management and reflection process.
@@ IV-D The Flexible Reflection Process
where $type$ represents the type of current task, $num_{obj}$ and $num_{inter}$ respectively represents the number of target objects and the number of interactions that need to be performed within the task. Among all episodes $ep_{total}$, the number of episodes for in-depth reflection is represented as:

### v2

@@ Abstract
Autonomous error correction is critical for domestic robots to achieve reliable execution of complex long-horizon tasks. Prior work has explored self-reflection in Large Language Models (LLMs) for task planning error correction; however, existing methods are constrained by inflexible self-reflection mechanisms that limit their effectiveness. Motivated by these limitations and inspired by human cognitive adaptation, we propose the Flexible Constructivism Reflection Framework (FCRF), a novel Mentor-Actor architecture that enables LLMs to perform flexible self-reflection based on task difficulty, while constructively integrating historical valuable experience with failure lessons. We evaluated FCRF on diverse domestic tasks through simulation in AlfWorld and physical deployment in the real-world environment. Experimental results demonstrate that FCRF significantly improves overall performance and self-reflection flexibility in complex long-horizon robotic tasks.
@@ III-B Self-Reflection Process in LLMs Planning
Discussing LLMs task planning process with self-reflection and memory module memmem, the self-reflection content for current unsuccessful planning trail $sr^{t}\in SR$ is generated by the LLM for self-reflection $M_{sr}$, while the reflection process can be described as $M_{sr}(sr^{t}|s^{t},mem^{t})$, meaning that the self-reflection content $sr^{t}$ of the current trail $t$ is generated by $M_{sr}$ based on the current state $s^{t}$, and persistent memory $mem^{t}$ lasts until the current trail including existing task trajectory.
The self-reflection generated by $M_{sr}$ would be stored in the agent memory memmem and passed to planner LLM in the next trail, to generate better action sequence $\{a^{t+1}_{1},\dots,a^{t+1}_{i}\}$ as the new output of the planning task. The sequence of actions is generated by the strategy function $\Phi(a^{t+1}_{1},\dots,a^{t+1}_{i}|G,s^{t+1},o^{t+1},sr^{t})$, meaning that new action sequence $\{a^{t+1}_{1},\dots,a^{t+1}_{i}\}$ from $step_{1}$ to $step_{i}$ is generated based on the task goal $G$, the current agent state $s^{t+1}$, the current observation of task environment $o^{t+1}$, and the generated self-reflection content $sr^{t}$ of last failed trail.
@@ IV-A LLM as Actor
where $A^{t}=\{a^{t}_{1},\dots,a^{t}_{i}\}$ denotes the current action sequence, $G$ represents the task goal, $o^{t}$ represents the current environmental observations during trail $t$, $SR^{t-1}=\{sr^{t-k},\dots,sr^{t-1}\}$ represents the set of self-reflection among the past $k$ trials to consider about, and $\theta$ represents the parameters of $M_{a}$. The action sequence generated by Actor LLM $M_{a}$ would be stored in the memory module memmem for the following self-reflection process of the task, until the current task is successfully completed.
@@ IV-B LLM as Constructivism Self-Reflection Mentor
The constructivism self-reflection process of our framework is guided by an LLM prompted as a $Mentor$, represented as $M_{m}$, which contains three sub-modules $M_{exp}$, $M_{lesson}$ and $M_{cons}$. The reflection process is divided into three parts: the summary of valuable experience, the summary of failure lessons, and the comprehensive construction of the two above parts. The final generated reflection content will be added to the memory module memmem and can be called by the Actor LLM $M_{a}$ in subsequent trails for planning.
@@
After completing the valuable experience summary process and the failure lesson summary process, the constructor submodule $M_{cons}$ will execute the comprehensive construction process to comprehensively construct the successful and failed experience, integrate the final reflection result in the form of an improved plan. During the construction process, $M_{cons}$ integrates the summarized valuable experience $Exp^{t}$ with the newly injected $Lesson^{t}$, unifying them into a sequence of actions, which forms the new constructive plan to be executed in the next trail. The reflected new constructive plan will be stored in the memmem module and guide the planning process of $M_{actor}$ in the next trail. The comprehensive construction process can be described as:
@@ IV-C Memory Management
During the task planning process, it is important for LLMs to maintain context memory. However, if all memory contents are included in prompts, it will lead to redundant prompts, increasing the memory burden of LLMs, even affect their performance. Motivated by the long-short-term memory mechanism of previous work Reflexion, we design the memory module memmem for comprehensive memory management, which is equivalent to an external buffer of LLMs. The memmem is divided into trajectory management module $Traj^{t}$ and reflection management module $Refl^{t}$, comprehensively manages the memory of task trajectories and reflection contents. The $Traj^{t}$ module is a short-term memory module storing the task trajectory $\tau^{t}$ of the current trail $t$. The $Refl^{t}$ module is a long-term memory module storing all failure self-reflection contents up to trail $t$, which can be represented as $\{sr^{1},\dots,sr^{t}\}$. The reflection contents are stored for targeted reading and calling by the Mentor LLM $M_{m}$ in the following trails. Our memory mechanism memmem relieves the memory pressure of LLMs, improving the efficiency of memory management and reflection process.
@@ IV-D The Flexible Reflection Process
where typetype represents the type of current task, $num_{obj}$ and $num_{inter}$ respectively represents the number of target objects and the number of interactions that need to be performed within the task. Among all episodes $ep_{total}$, the number of episodes for in-depth reflection is represented as: